OpenStudy (anonymous):

What is the solution to the equation 9^(x + 1) = 27 ?

OpenStudy (mertsj):

\[3^2^{(x+1)}=3^3\] \[3^{2x+2}=3^3\]

OpenStudy (mertsj):

Can you get it from there?

OpenStudy (anonymous):

Hint: log base 3 on both sides. :-)